A ring of mass (2π) kg and of radius 0.25 m is making 300 rpm about an axis through its perpendicular to its plane. The tension in Newton developed in ring is approximately(a)50 (b)100 (C)175 (d)250

Grade:11

1 Answers

Abhinav
5 years ago
Ans will be 250 as T=mw^2r/2π
Given m=2π
W^2=300*2π/60=[10π]^2=1000
T=2π*1000*0.25/2π
=1000*.25=250N
May this help you in making this question.This is hints for making question not the solution